# Efficient Nearest Neighbor Search Using Dynamic Programming

### v4

@@ Abstract
To address this, we propose a novel dynamic programming technique that precomputes a Directed Acyclic Graph (DAG) to encode the proximity structure between data points. More specifically, the DAG captures how the proximity structure evolves during the incremental construction of the Voronoi diagram of the data points. Experimental results demonstrate that our method achieves a 1x-10x speedup.
Additionally, our algorithm offers several valuable features. For instance, it naturally supports an $O(k\log n)$ algorithm for farthest point sampling, where $k$ is the desired number of sample points. Moreover, density peak clustering, which involves finding the nearest point among the top K points, is typically considered to have a time complexity of $O(n^{2})$. With our algorithm, this can be reduced to $O(n\log n)$. We believe this work will inspire further research on the NNS problem.
@@
Nearest neighbor search, farthest point sampling, density peak clustering, delaunay, voronoi diragram.
@@ I Introduction
  • Based on our method, we developed a farthest point sampling algorithm with an average complexity of $O(k\log n)$ in 3D, where $n$ represents the total number of points and $k$ denotes the number of samples.
@@ II Related Work
II-A Exact Nearest Neighbor Search
@@
Classic structures for space partitioning include KD Trees and quadtrees. A KD Tree [7] is a spatial partitioning structure that recursively divides the dataset along one of the $k$ coordinate system axes, typically alternating between them; for instance, with $k=2$, splits occur alternately on the x- and y-axes. This splitting process continues until a maximum depth is reached or until fewer than a specified number of points remain in a cell. KD Trees can be extended to include objects beyond points, such as triangles or line segments. Determining the optimal splitting plane can be challenging in these cases; however, for points (or disks) in two dimensions, one can simply sort the points along each dimension and split at the median. The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time.
@@
An octree is another spatial partitioning structure that specifically divides 3D space into eight cubes. This is typically done recursively, with each subdivision occurring when the data within a region meets certain criteria, such as exceeding a maximum number of points or objects. Each internal node in the octree corresponds to a cube region of space, further subdivided into eight equal-sized cubes represented by its children. The construction of an octree can also be completed in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time as well.
@@
The R Tree [8] is a classic structure for partitioning data. It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits, and all leaf nodes are at the same level. The R Tree can be built in $O(n\log n)$ time and allows for the identification of nearest neighbors in $O(\log n)$ time.
@@
Grids and linear search methods are also traditional approaches for nearest-neighbor searches. The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking agents within each cell. During a query, the search includes the cell containing the query points and the adjacent cells. The construction time for this method is $O(n+m)$, where $m$ represents the number of cells in the grid. While the worst-case query time is $O(n)$, practical performance greatly depends on the distribution of agents and the size of the cells. Conversely, linear search is suitable when the number of query points is very small. This method finds the nearest points by calculating the distance from each target point to the query point.
II-B Approximate Nearest Neighbor Search
Graph-based algorithms are highly effective for approximate nearest neighbor (ANN) search due to their balance of accuracy and computational efficiency [11, 12, 13, 14]. Typically, the process begins with constructing connections within a given point cloud, followed by iterative traversal of the graph starting from an entry point, which may be selected randomly or determined by a separate algorithm. At each step, the algorithm selects the closest neighboring node as the next base node, iteratively refining the current solution. The search terminates once a predefined stopping condition, such as a set number of distance evaluations, is met.
One well-known graph structure used for search is the Navigable Small World (NSW, also known as Metricized Small World, MSW) [15, 14, 12]. NSW constructs the graph by inserting elements in random order. Each element is then bidirectionally connected to its $M$ nearest neighbors among the previously inserted nodes. To improve search efficiency, many algorithms adopt a hierarchical strategy by creating multiple layers of the graph, which helps narrow the search space and accelerate the process.
Our method adopts a similar graph-based strategy, emphasizing efficient connectivity and traversal. Unlike traditional graph-based approximate nearest neighbor search, our method precisely identifies the nearest points with strict guarantees of correctness. Furthermore, although we do not explicitly use hierarchical structures, our integrated construction and traversal strategies achieve comparable benefits, such as efficient narrowing of the search space and improved accuracy.
II-C Voronoi Diagram & Delaunay
@@
Suppose we have a set of generators $\mathcal{S}=\{s_{i}\}_{i=1}^{n}$ in a given domain $\Omega$ equipped with a metric function $\mathcal{D}$. The Voronoi diagram partitions $\Omega$ into regions such that the generator $s_{i}$ dominates a region defined as
$\text{Cell}(s_{i})=\{x\in\Omega~{}|~{}\mathcal{D}(s_{i},x)\leq\mathcal{D}(s_{j},x),\,\forall j\neq i\}.$
The definition of Voronoi diagrams may vary based on domains, metrics, and types of generators. The most prevalent version assumes that $\Omega$ represents Euclidean spaces and that the generators are exclusively points. As a fundamental tool, Voronoi diagrams have found widespread applications in computer graphics [3, 1, 16].
@@
The Delaunay triangulation is the dual structure of the Voronoi diagram, representing the adjacency relationships between the Voronoi cells. Currently, Voronoi diagrams are typically computed by first calculating the Delaunay triangulation and then deriving the dual structure. The Watson algorithm [17] is a classic method for computing the Delaunay triangulation. It is based on the empty circle property and works by incrementally inserting points and flipping triangles that do not satisfy this property. Common tools for computing Delaunay triangulations include TetGen [18] and CGAL [19], which can achieve an average computational complexity of $O(n\log n)$ with a worst-case complexity of $O(n^{2})$.
Voronoi diagrams can also be applied to nearest neighbor searches. By definition, if the closest point to a query point $q$ in $\mathbf{P}$ is $p_{k}$, then $q$ lies within the Voronoi cell of $p_{k}$. This property makes Voronoi diagrams highly relevant for solving nearest neighbor problems. However, there is currently no efficient algorithm for identifying the specific Voronoi cell containing $q$. This restricts the practical use of Voronoi diagrams in nearest neighbor searches.
II-D Farthest Point Sampling
Farthest point sampling (FPS) is a technique widely employed in computational geometry and computer graphics applications. It is designed to iteratively select points from a dataset such that each newly selected point is maximally distant from the previously chosen points. This method can be formally expressed as follows:
$p_{i+1}=\arg\max_{p\in P}\left(\min_{p_{j}\in S}d(p,p_{j})\right),$ (2)
where:
  • $P$ represents the set of points in the point cloud,
  • $S$ denotes the subset of points that have already been selected,
  • $d(p,p_{j})$ is the distance between points $p$ and $p_{j}$,
  • $p_{i+1}$ is the next point selected such that its minimum distance to any point in $S$ is maximized.
This sampling strategy ensures that the selected points are well-distributed, making it effective for approximating shapes, surfaces, or distributions. FPS is particularly useful in applications requiring a representative subset of data points, such as mesh generation, surface simplification, or cluster initialization.
The time complexity of a typical FPS method [20, 21] is $O(kn)$, where $n$ is the total number of points in the dataset, and $k$ denotes the number of points to be sampled. The latest FPS method QuickFPS [22] organizes large-scale point clouds into multiple bins. By employing two mechanisms—merge computation and implicit computation on these bins—it significantly reduces external memory access and computation costs, thereby improving the speed of farthest point sampling.
@@ II-D Farthest Point Sampling
Figure 1: (a) The Voronoi diagram of $\{p_{i}\}_{i=1}^{20}$, where the query point $q$ (colored in red) belongs to the Voronoi cell colored in yellow. (b) If the addition of the three sites (whose cells are colored in green) does not change the yellow cell, $\Phi_{m}(q)$ remains unchanged when $m$ is increased from 20 to 23. (b) When the 24-th site $p_{24}$ is added (see the cyan cell), the yellow cell diminishes; yet $\Phi_{24}(q)=\Phi_{23}(q)$ if $q$ is nearer to the old site. (d) $\Phi_{24}(q)=p_{24}$ if $q$ is nearer to the newly added site $p_{24}$.
@@ III Methodology
$\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (3)
@@ IV-A Query Table Construction
Figure 2: We consider an incremental construction process of the Voronoi diagram of $\{p_{i}\}_{i=1}^{n}$. During the process, we append $p_{j}$ to $p_{i}$’s Query List if the addition of $p_{j}$ changes $p_{i}$’s cell. The rightmost figure shows the query process.
Since the Delaunay is the dual of the Voronoi diagram and encodes the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$. Algorithm 1 provides the pseudocode for the corresponding construction process.
@@ IV-B Nearest Neighbor Query
IV-C Farthest Point Sampling
Sections IV-A and  IV-B describe the algorithm for constructing the Query Table and performing nearest neighbor searches. The algorithm assumes a specific order of the point cloud and incrementally constructs the Voronoi / Delaunay based on this order. When the point cloud is unordered or there is no need to compute the nearest point to $q$ among the first $n$ points, we can sort the points in a specific manner to enhance the efficiency of the query phase.
Intuitively, the more uniformly the inserted point cloud is distributed in space, the faster the query speed. This is because it facilitates a transition from large to small spatial jumps during the query process. Farthest point sampling, a commonly used downsampling method, aims to distribute the sampled points as uniformly as possible across the point cloud. Drawing inspiration from this, we set the number of sampled points equal to the total number of points in the cloud and use the sampling order to incrementally construct the Voronoi diagram. Notably, by integrating this approach into our construction process, we developed a farthest point sampling strategy with reduced complexity, leading to faster sampling times.
@@ IV-C Farthest Point Sampling
As an example, consider a state from the incremental construction of the Voronoi diagram, as shown in Figure 3(a). The black points represent the points that have already been inserted, while the gray points denote the remaining points waiting to be inserted. When a new point is inserted based on the farthest point sampling strategy, as illustrated in Figure 3(b), the red point represents the newly sampled point from the gray points, and the green area corresponds to its Voronoi cell. A key observation is that only the distances of the remaining points located within the cells adjacent to the green cell (specifically, those within these cells prior to the insertion of the red point) need to be updated. Since the white Voronoi cell does not change during the insertion of new points, this implies that the nearest neighbors of the uninserted points within it remain unchanged. This significantly reduces the computational cost of sampling. The remaining task involves two parts: dynamically tracking the remaining points contained within the cells of the already inserted points during the incremental construction of the Voronoi diagram, and dynamically obtaining the next sampling point. The first part can be accomplished efficiently due to the local properties of Voronoi cells, while for the second part, we utilize a segment tree structure in our implementation.

### v5

@@ Abstract
To address this, we propose a novel dynamic programming technique that precomputes a Directed Acyclic Graph (DAG) to encode the proximity structure between data points. More specifically, the DAG captures how the proximity structure evolves during the incremental construction of the Voronoi diagram of the data points. Experimental results demonstrate that our method achieves a 1-10x speedup.
Additionally, our algorithm demonstrates significant practical value across diverse applications. We validated its effectiveness through extensive testing in four key applications: Point-to-Mesh Distance Queries, Iterative Closest Point (ICP) Registration, Density Peak Clustering, and Point-to-Segments Distance Queries. A particularly notable feature of our approach is its unique ability to efficiently identify the nearest neighbor among the first $k$ points in the point cloud—a capability that enables substantial acceleration in low-dimensional applications like Density Peak Clustering. As a natural extension of our incremental construction process, our method can also be readily adapted for farthest point sampling tasks. These experimental results across multiple domains underscore the broad applicability and practical importance of our approach.
@@
Nearest neighbor search, Delaunay triangulation, Voronoi diragram, farthest point sampling, density peak clustering,
@@ I Introduction
  • We demonstrate our algorithm’s practical effectiveness through comprehensive evaluation on four key applications: Point-to-Mesh Distance Queries, Iterative Closest Point Registration, Density Peak Clustering, and Point-to-Segments Distance Queries, showcasing its broad applicability in computational geometry scenarios.
@@ II Related Work
II-A Nearest Neighbor Search
@@
Classic structures for space partitioning include KD Trees and quadtrees. A KD Tree [7] is a spatial partitioning structure that recursively divides the dataset along one of the $k$ coordinate system axes, typically alternating between them; for instance, with $k=2$, splits occur alternately on the x- and y-axes. This splitting process continues until a maximum depth is reached or until fewer than a specified number of points remain in a cell. KD Trees can be extended to include objects beyond points, such as triangles or line segments. Determining the optimal splitting plane can be challenging in these cases; however, for points (or disks) in two dimensions, one can simply sort the points along each dimension and split at the median. The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time on average [11].
@@
An octree is another spatial partitioning structure that specifically divides 3D space into eight cubes. This is typically done recursively, with each subdivision occurring when the data within a region meets certain criteria, such as exceeding a maximum number of points or objects. Each internal node in the octree corresponds to a cube region of space, further subdivided into eight equal-sized cubes represented by its children. The computation cost for locating a leaf node in an octree is $O(depth)\approx O(\log n)$, where depth represents the average depth of the tree [12]. However, nearest neighbor queries typically require additional traversal operations beyond simple leaf node access, potentially increasing the practical query time by either constant factors or additional complexity terms.
@@
The R Tree [8] is a classic structure for partitioning data. It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits, and all leaf nodes are at the same level. The R*-tree [13] is an enhanced variant that incorporates a combined optimization of area, margin, and overlap of enclosing rectangles. It introduces forced reinsert to dynamically reorganize the structure and prevent unnecessary splits, demonstrating superior performance over existing R-tree variants across different data distributions and query types.
@@
Grids and linear search methods are also traditional approaches for nearest-neighbor searches. The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking agents within each cell. During query operations, the algorithm examines the cell containing the query point and its adjacent cells. However, this approach exhibits significant performance degradation in certain scenarios—particularly when data points are distributed on a spherical surface with the query point at the center. In such cases, the algorithm must traverse multiple cells in all directions, essentially performing a comprehensive dataset scan. Linear search, conversely, offers a straightforward solution by calculating distances between the query point and each target point, making it suitable only when dealing with very few query operations or small datasets.
Additionally, sketching techniques have been proposed for similarity search acceleration. Mic et al. [14] introduced methods for selecting optimal sketches that transform objects into bit-strings. Their sketch construction often relies on reference points, or pivots, to partition the data space. While their approach focuses on Hamming embedding for approximate similarity filtering, our method constructs a ‘Query Table’ based on the incremental insertion of points, which begins the search from a fixed initial point. Our goal is the exact nearest neighbor search in metric spaces through Voronoi diagram construction.
II-B Voronoi Diagram & Delaunay
@@
Given a set of generators $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$ in a domain $\Omega$ equipped with a metric function $\mathcal{D}$, the Voronoi diagram partitions $\Omega$ into regions where each generator $p_{i}$ dominates a cell defined as:
$\text{Cell}(p_{i})=\{x\in\Omega~|~\mathcal{D}(p_{i},x)\leq\mathcal{D}(p_{j},x),\,\forall j\neq i\}.$ (2)
While Voronoi diagrams can be generalized across different domains, metrics, and generator types, the most commonly studied variant assumes Euclidean spaces with point generators. As fundamental geometric structures, Voronoi diagrams have found extensive applications in computer graphics [3, 1, 15].
@@
The Delaunay triangulation serves as the geometric dual of the Voronoi diagram, encoding the adjacency relationships between Voronoi cells. In 2D, the Delaunay triangulation consists of triangles whose vertices correspond to the Voronoi generators, while in 3D, it comprises tetrahedra with the same vertex correspondence. Figure 1 illustrates this duality in the 2D setting, where the Delaunay triangulation is depicted with black dashed lines and the Voronoi diagram with gray solid lines. This organizational structure makes it straightforward to obtain the adjacent sites of any given vertex, with many computational geometry libraries providing corresponding interfaces for such operations. Currently, Voronoi diagrams are typically computed by first calculating the Delaunay triangulation and then deriving the dual structure. The Bowyer–Watson algorithm [16] is a classic method for computing the Delaunay triangulation. It is based on the empty circle property and works by incrementally inserting points and flipping triangles that do not satisfy this property. Common tools for computing Delaunay triangulations include TetGen [17] and CGAL [18], which can achieve an average computational complexity of $O(n\log n)$ with a worst-case complexity of $O(n^{2})$ [19].
@@
Figure 1: Example of Voronoi diagrams and Delaunay triangulations in 2D. The gray solid lines represent the Voronoi diagram partitioning the plane based on proximity to point generators (black dots), while the black dashed lines show the corresponding Delaunay triangulation that encodes the connectivity relationships between neighboring Voronoi cells.
The geometric properties of Voronoi diagrams also enable nearest neighbor queries. By construction, if the closest point to a query location $q$ among generators $\mathbf{P}$ is $p_{k}$, then $q$ necessarily lies within $\text{Cell}(p_{k})$. However, the practical utility of this property is limited by the absence of efficient algorithms for point location within Voronoi cells, presenting an ongoing challenge for large-scale nearest neighbor applications.
[…]
Figure 2: (a) The Voronoi diagram of $\{p_{i}\}_{i=1}^{20}$, where the query point $q$ (colored in red) belongs to the Voronoi cell colored in yellow. (b) If the addition of the three sites (whose cells are colored in green) does not change the yellow cell, $\Phi_{m}(q)$ remains unchanged when $m$ is increased from 20 to 23. (c) When the 24-th site $p_{24}$ is added (see the cyan cell), the yellow cell diminishes; yet $\Phi_{24}(q)=\Phi_{23}(q)$ if $q$ is nearer to the old site. (d) $\Phi_{24}(q)=p_{24}$ if $q$ is nearer to the newly added site $p_{24}$.
@@ III Methodology
$\hskip-8.53581pt\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (3)
@@ IV-A Query Table Construction
Figure 3: We consider an incremental construction process for the Voronoi diagram and its corresponding Delaunay triangulation of $\{p_{i}\}_{i=1}^{n}$. During this process, we append $p_{j}$ to $p_{i}$’s Query List if the insertion of $p_{j}$ modifies $p_{i}$’s Voronoi cell. In other words, within the Delaunay triangulation, $p_{j}$ and $p_{i}$ are adjacent neighbors. The rightmost figure illustrates this query process.
Since the Delaunay is the dual of the Voronoi diagram and encodes the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$. Figure 3 illustrates the synchronized incremental construction process of both Voronoi diagrams and their dual Delaunay triangulations. As Delaunay triangulations explicitly encode structural relationships between points, obtaining neighboring points becomes straightforward, with most computational geometry libraries providing efficient interfaces for this purpose. Algorithm 1 provides the pseudocode for the corresponding construction process.
@@ IV-B Nearest Neighbor Query
[…]
Figure 4: Coupling of Farthest Point sorting with incremental Voronoi Construction. (a): Uninserted points (gray) are distributed within the Voronoi cells of the inserted points (black). (b): When a new point is inserted, only the distance of the uninserted points within the green and yellow cells may need to be updated.
IV-C Insertion Reordering Based on Farthest-Point Sorting
Sections IV-A and  IV-B describe the algorithm for constructing the Query Table and performing nearest neighbor searches. The algorithm assumes a specific order of the point cloud and incrementally constructs the Voronoi / Delaunay based on this order. When the point cloud is unordered, we can sort the points in a specific manner to enhance the efficiency of either the preprocessing or query phase. Intuitively, the more uniformly the inserted point cloud is distributed in space, the faster the query speed, as it facilitates a transition from large to small spatial jumps during the query process.
Farthest point sampling (FPS) is a technique that iteratively selects points from a dataset such that each newly selected point is maximally distant from the previously chosen points. This process can be formalized as:
$p_{i+1}=\arg\max_{p\in P}\left(\min_{p_{j}\in S}d(p,p_{j})\right),$ (6)
where $P$ represents the point cloud, $S$ denotes the already selected points, $d(p,p_{j})$ is the distance between points, and $p_{i+1}$ is the next selected point that maximizes its minimum distance to any point in $S$. This strategy ensures well-distributed point selection, making it effective for shape approximation and surface representation.
Drawing inspiration from this, we set the number of sampled points equal to the total number of points in the cloud and use the sampling order to incrementally construct the Delaunay triangulation. Notably, by integrating this approach into our construction process, we developed a farthest point sampling strategy with reduced complexity, leading to faster sampling times.
@@ IV-C Insertion Reordering Based on Farthest-Point Sorting
As an example, consider a state from the incremental construction of the Voronoi diagram, as shown in Figure 4(a). The black points represent the points that have already been inserted, while the gray points denote the remaining points waiting to be inserted. When a new point is inserted based on the farthest point sampling strategy, as illustrated in Figure 4(b), the red point represents the newly sampled point from the gray points, and the green area corresponds to its Voronoi cell. A key observation is that only the distances of the remaining points located within the cells adjacent to the green cell (specifically, those within these cells prior to the insertion of the red point) need to be updated. Since the white Voronoi cell does not change during the insertion of new points, this implies that the nearest neighbors of the uninserted points within it remain unchanged. This significantly reduces the computational cost of sampling.
The remaining task involves two parts: dynamically tracking the remaining points contained within the cells of the already inserted points during the incremental construction of the Voronoi diagram, and dynamically obtaining the next sampling point. For the first task, we dynamically maintain a vector for each inserted point, storing the remaining points that are nearest to that point. We update these vectors dynamically during the incremental Delaunay construction process. For the second task, a segment tree [21] provides an excellent data structure that supports both single-point value updates within the interval and maximum value queries within same interval, both in logarithmic time. Therefore, during incremental Delaunay construction, we obtain all remaining points requiring distance updates through vector updates, update them in the segment tree (with inserted points set to negative infinity), and then leverage the segment tree’s logarithmic-time maximum query capability to retrieve the next farthest sampling point.
IV-D Two Delaunay Construction Strategies
For the construction of Delaunay triangulations, numerous algorithms exist; we employ the Bowyer–Watson algorithm [16], which performs incremental Delaunay construction through sequential point insertion. When inserting a new point, the algorithm first locates the Delaunay tetrahedron containing the point’s spatial location, then performs edge flipping operations based on the empty sphere property until the Delaunay criteria are satisfied. In our framework, different insertion orders for the Delaunay construction can achieve either faster construction speed or improved query performance. We introduce two strategies adopted in this work:
Prioritizing Construction Speed. One time-consuming operation during Delaunay construction is locating the tetrahedron containing the newly inserted point. To accelerate Delaunay construction, a common approach involves pre-sorting the input point cloud using space-filling curves before performing incremental construction. This strategy ensures that the point insertion order exhibits spatial adjacency properties, enabling newly inserted points to quickly propagate from the previous insertion location to locate the tetrahedron containing the new point. Additionally, this approach provides excellent spatial locality during the construction process.
We adopt the point cloud sorting strategy from CGAL [22], specifically the spatial_sort function, which first organizes the point set into random buckets of increasing sizes, with Hilbert sorting applied only within each bucket. Our experiments demonstrate that this strategy achieves fast Delaunay construction speed with minimal impact on query performance.
Prioritizing Query Performance. Farthest Point Sorting ensures that inserted points are distributed as uniformly as possible throughout the full point cloud, which intuitively leads to faster query speeds—a hypothesis validated by our experiments. The farthest point sampling strategy has been detailed in Section IV-C. Here, we primarily focus on how our algorithm can be integrated to accelerate the Delaunay construction process under the farthest point sampling ordering.
Our algorithm possesses an important property: as illustrated in the Figure 3 (b,c), before inserting the 5th point, the query table for the first four points has already been constructed (shown in the bottom portion of the Figure 3 (b) ), enabling queries for the nearest point among the first 4 points. Consequently, when inserting point $p_{k}$, we can already utilize Algorithm 2 to query the nearest point to $p_{k}$ within $\{p_{i}\}_{i=1}^{k-1}$.
As previously mentioned, one time-consuming operation during Delaunay construction is locating the tetrahedron containing the newly inserted point. Therefore, when inserting the $k$-th point, we can quickly determine the nearest point to $p_{k}$ from $\{p_{i}\}_{i=1}^{k-1}$. This nearest point then serves as the starting location for efficiently locating the tetrahedron containing $p_{k}$, significantly accelerating the Delaunay construction process.
IV-E Time Complexity Analysis
Preprocessing. Our preprocessing phase involves the incremental construction of the Delaunay triangulation, during which we store the requisite information for subsequent queries. Consequently, the preprocessing time complexity aligns with that of Delaunay triangulation construction.
On average, the time complexity for constructing a Delaunay triangulation in 3D is $O(n\log n)$ [23], while the worst-case time complexity is $O(n^{2})$ [19]. However, this worst-case scenario is rarely encountered in typical construction processes.
Querying. Due to the difficulty in evaluating the impact of point cloud distribution on Delaunay construction, providing a rigorous theoretical proof for average-case complexity remains challenging. Therefore, we present an experimental average complexity based on empirical observations, followed by a theoretical worst-case complexity analysis. The experimental section validates the correctness of our experimental average complexity analysis.
Experimental Average Complexity
When querying the nearest neighbor for point $q$, the complexity of the query phase depends on the number of times the Voronoi cell containing $q$ undergoes shrinkage during the incremental construction process. Each time the Voronoi cell containing $q$ shrinks due to the insertion of a new point, a comparison operation is required to potentially update the nearest neighbor.
More specifically, assume $\Phi_{m}(q)$ represents the current nearest neighbor of $q$ among the first $m$ points, and a new point $p_{m+1}$ is inserted. Based on empirical observations, on average, the Voronoi cell of $p_{m+1}$ is adjacent to a constant number of existing cells, denoted as $c$ [24]. Consequently, the probability that $p_{m+1}$ is adjacent to $\Phi_{m}(q)$ in the Delaunay triangulation is approximately $c/m$, which corresponds to the likelihood of requiring a comparison operation.
Thus, the average time complexity of the nearest neighbor query phase is:
$O\left(\sum_{i=1}^{n-1}\frac{c}{i}\right)=c\times O(\log n)=O(\log n).$ (7)
Theoretical Worst-Case Complexity
The theoretical worst-case complexity of our algorithm is $O(n)$, as it relies on pairwise point comparisons and eliminates at least one invalid candidate with each comparison. A typical example is when all points in 3D space lie on a straight line, ordered sequentially as $p_{1},p_{2},\ldots,p_{n}$, with Delaunay construction following this insertion order. In this degenerate configuration, each point $p_{k}$ encroaches only upon $p_{k-1}$, creating a chain-like structure that forces linear traversal during nearest neighbor queries, thereby degrading the algorithm to $O(n)$ complexity.